OpenStudy (anonymous):

find the variation of constant and an equation of variation where y varies as x and y =36 when x=9.

OpenStudy (amistre64):

we have the relationship: y = x 36 = 9 ; this is jibberuish as is, to make any sence of it we need to introduce a "constant of variation" (k) 36 = 9k ; but what should k be for this to make sense? k = 36/9 = 4 36 = 9(4) y = 4x